OpenStudy (anonymous):

6x^2+13x-5 plz help!

jimthompson5910 (jim_thompson5910):

Multiply the first coefficient and the last number 6 and -5 to get 6*(-5) = -30 Now find two numbers that multiply to -30 and add to 13 at the same time.

OpenStudy (anonymous):

well 5 times 6 = 30

OpenStudy (anonymous):

but idk how to add it up and = 13

jimthompson5910 (jim_thompson5910):

but 5+6 = 11, so that's out

jimthompson5910 (jim_thompson5910):

what else multiplies to 30?

jimthompson5910 (jim_thompson5910):

other than 5 and 6

OpenStudy (anonymous):

hmm 15 times 2

OpenStudy (anonymous):

3 times 10

jimthompson5910 (jim_thompson5910):

good, so which one is negative?

jimthompson5910 (jim_thompson5910):

your first choice of 15 and 2 is the right one (sort of), but one of them is negative

OpenStudy (anonymous):

o hmm -2?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 15 and -2

OpenStudy (anonymous):

= -30

OpenStudy (anonymous):

now when u add them both together u get -13 right?

jimthompson5910 (jim_thompson5910):

more like +13, but you have the right idea

OpenStudy (anonymous):

o my bad yea so it would be (x+15)(x-2)?

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (anonymous):

ooohh

jimthompson5910 (jim_thompson5910):

because that first coefficient is NOT 1, we can't use the trick I showed you before

jimthompson5910 (jim_thompson5910):

but the two numbers we just found are still important

OpenStudy (anonymous):

oook so what should i do now?

jimthompson5910 (jim_thompson5910):

break up 13x into 15x - 2x to get 6x^2+15x - 2x -5 Then factor this by grouping

jimthompson5910 (jim_thompson5910):

Do you know how to do that ?

OpenStudy (anonymous):

hmmm not really can u show me?

jimthompson5910 (jim_thompson5910):

6x^2+15x - 2x -5 (6x^2+15x) + (-2x -5) .... pair off the terms in parenthesis 3x(2x+5) - 1(2x + 5) ... factor out 3x from the first group and -1 from the second (3x - 1)(2x + 5) ... factor out the GCF 2x+5 So 6x^2+15x - 2x -5 factors to (3x - 1)(2x + 5) By extension, this means 6x^2+13x-5 factors to (3x - 1)(2x + 5)

OpenStudy (anonymous):

so would it be 6x^2+13x-5=(3x-1)(2x+5)?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you sooo much

jimthompson5910 (jim_thompson5910):

yw